A disk of mass M is spinning freely at 4.49 rad/s when a second identical disk, initially not spinning, is dropped onto it so that their axes coincide. In a short time the two disks are corotating.

a) What is the angular speed of the new system (in rad/s)?

b)  If a third such disk is dropped on the first two, find the final angular speed of the system (in rad/s).
Answer #1

given
wo = 4.49 rad/s

a) let I is the moment of inertia of the disk.

let w is the final angular speed.

Apply conservation of anghular momentum

Lf = Li

(I + I)*w = I*wo

w = wo/2

= 4.49/2

= 2.245 rad/s <<<<<<<<---------------Answer

b) let w' is the final angular speed

Lf = Li

(I + I + I)*w' = 2*I*w

w' = 2*w/3

= 2*2.245/3

= 1.50 rad/s <<<<<<<<---------------Answer